Required information E11-4 (Algo) Reporting Stockholders Equity LO11-1, 11-2, 11-3 [The following information applies to theRequired information E11-4 (Algo) Reporting Stockholders Equity LO11-1, 11-2, 11-3 [The following information applies to theRequired information E11-4 (Algo) Reporting Stockholders Equity LO11-1, 11-2, 11-3 [The following information applies to theRequired information E11-4 (Algo) Reporting Stockholders Equity LO11-1, 11-2, 11-3 [The following information applies to theE11-10 (Algo) Computing and Interpreting Earnings per Share LO11-2 Below is select information from DC United Companys incom
Answer #1

1) Additional paid in capital = (109000*21-1090000) = 1199000

2) Beginning retained earnings = 790000+620000-1180000 = 230000

3) Treasury Stock = 109000-77000 = 32000 Share

4) earning per share = 1180000/77000 = 15.32 per share

5) earning per share = 79500/138500 = 0.57 per share